Given a 550 L sample of neon gas that has a temperature of 23 degrees c and a pressure of 765 mmh, determine how many moles are present. how many grams is this?
1
Expert's answer
2019-04-09T07:54:57-0400



Task:

Given a 550 L sample of neon gas that has a temperature of 23 degrees and a pressure of 765 mmHg. Determine how many

moles are present? How many grams is this?


Solution:

According to ideal gas equation: PV = nRT

Pressure = P = 765 mmHg = 1 atm.

Volume = V = 550 L.

Temperature = T = 23°C = (23 + 273.15) = 296.15 K.

Gas constant = R = 0.08206 L*atm*mol-1*K-1.

Then,

n(Ne, g) = PV / RT = (1 atm * 550 L) / (0.08206 L*atm*mol-1*K-1 * 296.15 K) = 22.6318 mol.

n(Ne) = 22.6318 mol.

M(Ne) = 20.1797 g/ mol.

n(Ne) = m(Ne) / M(Ne);

m(Ne) = n(Ne) * M(Ne) = 22.6318 * 20.1797 = 456.703 g

m(Ne) = 456.703 g


Answer: n(Ne) = 22.632 mol; m(Ne) = 456.703 g.